If x+3,3x-4 are in AP. Find the value of x

Answer by richwmiller(17219)

Not enough info for just one solution
3x-4-(x+3)=2x-7
3x-4+2x-7=5x-11 next term
5x-11-(3x-4)=3x-4-(x+3)
x+3-(2x-7) = -x-4 previous term
d = 2n+1, x = n+4